OpenStudy (anonymous):

question on groups and symmetries 1) The dihedral group of order 12 is D6 = {e, r, r^2, r^3, r^4, r^5, s, rs, (r^2)s, (r^3)s, (r^4)s, (r^5)s } where |r| = 6, |s| = 2 and sr = (r^5)s a) Let H = and show that H is normal to D6( the sideway triangle symbol). b) Construct the Cayley table of D6/H. Is this group abelian? Is it cyclic?

OpenStudy (anonymous):

@nbouscal

OpenStudy (anonymous):

to show that H is normal to D6 we have to show that aH = Ha right?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

to show it is normal this is what i did

OpenStudy (anonymous):

(e)H = {e, r^2, r^4} = H(e) (r)H = {r, r^3, r^5} = H(r) (r^2)H = {r^2, r^4, e } = H(r^2) (r^3)H = {r^3, e} = H(r^3) (r^4)H = {r^4, r^2} = H(r^4) (r^5)H = {r^5, r^4, r^2} = H(r^5) and i did this for the s, rs, .... but could yoi please tell me if i did the above right and how on earth do i make a cayley table out of this

OpenStudy (anonymous):

@KingGeorge

OpenStudy (kinggeorge):

The above looks correct to me. That successfully shows that H is normal in \(D_6\). To construct the Cayley table, start with two things. First, write all your group elements across the top and the side (in the same order, and e is usually first). It should look something like the following. e r r^2 r^3 r^4 ................ e r r^2 r^3 r^4 r^5 s sr sr^2 sr^3 sr^4 sr^5 Make sure to continue the top line.

OpenStudy (kinggeorge):

Once you've drawn all that out, the entries in the table will be \(gh\), where \(g\) is the row label, and \(h\) is the column label.

OpenStudy (anonymous):

alright perfect i shall follow

OpenStudy (anonymous):

but is D6/H a significance ?

OpenStudy (anonymous):

@KingGeorge

OpenStudy (kinggeorge):

While making the Cayley table, it's irrelevant that H is normal to D6

OpenStudy (anonymous):

ok i can clearly see that this is abelian but how would i know its cyclic i see a pattern but how exactly would i word that this D6 group is cyclic?

OpenStudy (kinggeorge):

D6 should not be abelian.

OpenStudy (anonymous):

wth that is weird i am getting it is abelian :s

OpenStudy (anonymous):

alright i am going to be posting a picture

OpenStudy (anonymous):

@KingGeorge

OpenStudy (anonymous):

i stopped right there but whatever i pick i get it being abelian or did i draw the cayley table wrong

OpenStudy (kinggeorge):

I think you're drawing it a bit wrong. You seem to be switching between \(gh\) and \(hg\). It's important that you remain consistent throughout the entire table.

OpenStudy (anonymous):

hmm what do you mean by that :s wont it always be the same like

OpenStudy (anonymous):

like say i do (r^4)o(r^3) = r^7 = r right

OpenStudy (anonymous):

and by the table i get (r^3)o(r^4) = r^7 = r also :s

OpenStudy (kinggeorge):

That's all fine. However, when you start multiplying elements with s, things get a bit weird. For example, in the eighth row and second column, you have \(rsr=r^2s\), when it should be \(rsr=rr^5s=s\).

OpenStudy (anonymous):

o really :$ and thats by the condition they gave us whereever we see sr we replace it with r^5*s?

OpenStudy (kinggeorge):

Correct. \(sr=r^5s\).

OpenStudy (anonymous):

and how would i knwo it is cyclic?

OpenStudy (kinggeorge):

You would know it's cyclic if you can find some element \(g\) such that \(|g|=12\).

OpenStudy (anonymous):

well clearly e is no?

OpenStudy (kinggeorge):

e is definitely not.

OpenStudy (kinggeorge):

I think an easy way is to look at the diagonal of the table you've constructed. If every element in the diagonal is different, then your group is cyclic.

OpenStudy (kinggeorge):

Let me know what you've got. After you post something I can post the table I've made.

OpenStudy (anonymous):

alright i am still drawing the table i made it again ill finish half of it and just tell me if it is right

OpenStudy (kinggeorge):

Sounds good. Just post it, and I'll let you know.

OpenStudy (anonymous):

@KingGeorge

OpenStudy (kinggeorge):

That looks good so far. I would also recommend always writing elements in the form \(r^ns\) if you have both r's and s's in some element. For example, instead of writing \(sr^4\), write \(r^2s\). You don't have to edit what you've already written, or start over, just follow this for the rest of the matris.

OpenStudy (anonymous):

hmm i see what you mean

OpenStudy (anonymous):

and just to clarify s*s is just s right

OpenStudy (kinggeorge):

Also, I don't think what I said earlier with unique entries on the diagonal was correct. Just ignore that comment. \(s*s=s^2=e\)

OpenStudy (anonymous):

ooo yes truee truee and yeah i was going to ask you about that i was getting some similar terms

OpenStudy (kinggeorge):

I'll post the table I made now. It's a little bit different format, but you should be able to check against it.

OpenStudy (anonymous):

hmm alright thanks a lot @KingGeorge and we can see this is cyclic because

OpenStudy (anonymous):

ok you said they are suppose to have order of 12 right

OpenStudy (anonymous):

each row has a different element though

OpenStudy (kinggeorge):

There must be an element with order 12 for \(D_6\) to be cyclic. If you look though, you should find that \(D_6\) is not cyclic.

OpenStudy (anonymous):

hmm because there is no cycle occuring

OpenStudy (kinggeorge):

Precisely.